# Visualizing Thought: Conceptual Diagrams Enable Robust Planning in LMMs

### v1

@@ 2 Method
Models Claude 3.5 (Sonnet) Claude 3 (Opus) GPT-4o GPT-4 Turbo LLaMA 3.1 405B Gemini 1.5 Pro o1-preview o1-mini GPT4-o + Visual Thinking One Shot 57.6% 48.1% 28.3%* 23% 47.3% 16.8% - - - Zero Shot 54.8% 59.3% 35.5% 40.1% 62.6% 23.8% 97.8% 56.6% 90.2%
Table 1: Performance of different models in the Blocksworld domain. Baseline numbers taken from [54]
3 Related Work
Multimodal Chain-of-Thought for Reasoning. Several recent works have explored integrating visual representations into reasoning processes for LLMs and LMMs. [26] equips multimodal models with drawing tools to generate intermediate sketches composed of lines and shapes, but primarily addresses single-step problems with only a single intermediate visualization per problem instance. Similarly, [56] generates visual aids for spatial reasoning, providing a single refined visualization per problem, which typically is an approximate of a high-fidelity illustration rather than a conceptual diagrams representing object interactions. Concurrent work, [59], generates visual and textual intermediate steps for planning tasks by conditioning multi-modal models on externally provided initial visual representations that supplement textual problem descriptions. Their approach iteratively generates intermediate textual and visual states, but the reliance on externally supplied initial visualizations could limit generality and applicability to unseen domains.
Our approach differs in several key aspects. First, our method autonomously generates the first conceptual diagram directly from textual problem descriptions, without relying on external visual demonstrations or cues. Second, rather than producing a single visualization, our approach creates multiple chains of intermediate visual states, enabling parallel multi-hypothesis compositional reasoning through evolving diagrams. Third, our diagrams explicitly represent relationships between objects in a conceptual manner, rather than realistic or approximate visual depictions of the objects. Finally, our method operates in a zero-shot setting, capable of solving any problem expressible in PDDL format (very general, compatible with most planning problems) without prior visual examples or domain-specific visual guidance. These distinctions collectively enable a more generalizable and flexible form of diagrammatic reasoning, leading to significant performance gains.
Search and Verification Strategies for Enhanced LLM Reasoning. Recent methods improve reasoning in LLMs by structuring inference into explicit intermediate steps [58, 29] and employing search-based strategies over multiple reasoning paths [57, 61, 21, 11, 5]. For instance, Yao et al. [61] propose Tree-of-Thoughts (ToT), which extends CoT to explore a tree of reasoning paths, and Besta et al. [5] introduce Graph-of-Thoughts (GoT), which structures the reasoning process as a graph, enabling exploration and aggregation of intermediate reasoning steps. Other works utilize verification and refinement techniques, such as iterative self-reflection and feedback, to enhance reasoning accuracy [36, 43, 48]. Our approach builds upon these methods by integrating conceptual diagrams into intermediate reasoning steps. Moreover, as detailed in Section 2, we extend graph-based inference with beam search and depth-wise backtracking, significantly reducing the search space and improving performance on long-horizon planning tasks.
World Modeling and Planning with Language Models.. Recent research explores planning and reasoning using LLMs by implicitly or explicitly constructing world models from textual descriptions [27, 60, 3, 13, 14, 33, 55]. For instance, Huang et al. [27] show that LLMs implicitly form textual world models for simple planning tasks, while other works explicitly represent states and transitions for structured reasoning [60, 13, 33]. Recent approaches also introduce explicit state-tracking, modeling entity relationships and transitions to enhance reasoning [33, 55, 14]. Our work extends these approaches by enabling LMMs to autonomously construct and reason with diagrammatic world models. Specifically, our method generates conceptual diagrams directly from textual problem descriptions, defining a visual schema of objects and their relationships through abstract shapes, colors, and spatial relationships. Subsequent diagram updates visually simulate action sequences, explicitly tracking object states and interactions. Compared to purely textual world modeling, our diagrammatic approach provides clearer and more accurate representations of spatial and relational information, significantly improving performance on complex, multi-step planning tasks.
Domain GPT4o o1-mini o1-preview GPT4-o + Visual Thinking
Blocksworld (hard) 0% 10% 23.65% 78%
Floor Tiles 0% 6.7% 23.3% 36.7%
Parking 3.3% 6.7% 36.7% 50%
Tetris 0% 3.33% 26.67% 40%
Elevator 0% 3.33% 33.3% 46.67%
Barman 0% 0% 13.3% 33.3%
Table 2: Performance of select models on various domains.
@@ 4 Evaluation
To evaluate the proposed approach we have conducted experiments 7 different planning problem domains, including the popular Blocksworld [20] and Blocksworld Hard [54]; as well as 5 new problem domains contributed in this work, prepared by us using a popular repository of PDDL planning problem domains [1].
4.1 Implementation Details
We evaluated our method using GPT-4o as the LMM, and o1-preview [42] and o1-mini [41] as baselines, all accessed via the OpenAI API. Baseline models solved instances in PDDL format using prompts from [54]. Plan correctness was determined with VAL [25]. Experiments ran on a machine with 96 Intel Xeon 5220R CPUs (4GHz, 71.5 MiB L3 cache). On this machine, processing each depth level took approximately 1 minute. Consequently, simpler Blocksworld instances typically took around 5–8 minutes per instance, which is comparable to the runtime of the o1-preview model. Our code is provided in the supplementary and would be released upon acceptance.
Analysis Corrects Incorrects Incompletes Avg Depth of Corrects Max Depth of Corrects Min Depth of Corrects Avg Number of States Generated
@@ 4.1 Implementation Details
Floor Tiles 36.7% 23.3% 40% 15.88 25 10 244
Parking 50% 40% 10% 9.6 29 2 149.1
Tetris 40% 60% 0% 6.45 11 4 46.24
Elevator 46.67% 43.33% 10% 20.89 26 16 160.30
Barman 33.33% 40% 26.67% 24.0 27 22 210.54
@@
Table 3: Analysis of domains. ‘Corrects’: % instances with correct plans; ‘Incorrects’: % with incorrect plans; ‘Incompletes’: % terminated due to state budget; ‘Avg/Max/Min Depth of Corrects’: avg/max/min number of actions in correct plans; ‘Avg Number of States Generated’: avg states generated across all instance of the domain. ‘Corrects’ is our main metric.
Ablation Corrects Incorrects Incompletes Avg Depth of Corrects Max Depth of Corrects Avg Number of States Generated
GPT4-o + Visual Thinking 90.2% 6.4% 3.4% 10.07 28 38.89
No Diagram 58% 36% 6% 8.28 18 25.86
@@
Table 4: Ablation Study Results. ‘Corrects’ is our main metric. See Tab. 3 for columns notation.
4.2 Datasets
Evaluated Domains.
We evaluate our method on combinatorial planning problems [17] from the International Planning Competition (IPC) [2], expressed in PDDL format [38]. These problems are mostly NP-hard [20]. We evaluated out method on six domains, namely Blocksworld [20], and 5 other IPC domains: Floor Tile [23], Parking, Tetris [51], Elevator, and Barmen. For each domain, we generated instances using publicly available instance generators [1]. The Blocksworld domain involves stacking and unstacking blocks. We used the 500 simple instances in PlanBench [52] with 3-5 blocks and 30 additional hard instances with 10-20 blocks (same instance complexity as in [54]). The Floor Tile domain features robots painting tiles on a grid. We generated 30 PDDL instances for this domain with varying grid sizes (2-3 rows, 3-5 columns) and 1-2 robots. The Parking domain involves rearranging cars in curbs, with 30 instances using 4-5 curbs and 4-6 cars. The Tetris domain requires rearranging Tetris tiles on a grid, with 30 instances using $(4\times 4)$ or $(6\times 6)$ grids. Lastly, the Elevator domain simulates passenger transport in buildings, with 30 instances using 4-5 floors and 10-12 passengers. Refer to Fig. 4 a sequence of intermediate state diagrams in the correct chain for an example instance of a subset of evaluated domains.
Translating PDDL to Natural Language and Back.
Our method operates entirely over natural language descriptions. Thus, we first translate PDDL domains and instances into natural language. For domain translation, we use a manually engineered five-shot prompt containing translations of five different PDDL domains into natural language. For instance translation, we use a one-shot prompt containing a manually translated instance from the same domain. After solving the problem using our pipeline, we translate the resulting natural language action sequence back into PDDL using a one-shot prompt, which includes a random partial plan from the same domain along with its syntax-accurate translation to PDDL.
4.3 Results
Our main results are summarized in Tables 1 and 2. Table 1 compares the performance of our method (using GPT-4o as the LMM) on the popular, yet simpler Blocksworld planning domain  against leading closed- and open-source models at the time of writing, including the specialized reasoning models o1-preview and o1-mini. Most baseline results are taken from [54]. Our approach significantly outperforms the base GPT-4o model by approximately 55%, achieving 90.2% accuracy, and outperfoms all baselines, except for a slight disadvantage compared to the leading reasoning model, o1-preview (97.8%). We conjecture two reasons for o1-preview’s advantage: (i) the smaller problem size and shallow solution depth of simpler Blocksworld instances, and (ii) This planning domain is widely used, with many solved instances publicly available online, likely providing extensive exposure to o1 during its training.
This conjecture is supported by our results in Table 2, which evaluates performance on significantly larger and deeper Blocksworld (hard) instances, as well as five additional challenging domains derived from the IPC competition [2], where new harder instances are generated and solutions are not publicly available. In these more complex and unfamiliar domains, our method substantially outperforms all baselines, including o1-preview. For instance, in Blocksworld (hard), our method achieves 78% accuracy compared to only 23.65% by o1-preview. Similarly, our method consistently surpasses o1-preview across Floor Tiles, Parking, Tetris, Elevator, and Barman domains, demonstrating the robustness and generalizability of our structured visual reasoning approach.
Table 3 provides further insights into our method’s performance on these challenging domains. Despite significantly deeper solution paths (with average depths up to 24 actions per instance and up to 244 states generated per instance), our method successfully solves instances requiring as many as 36 sequential actions (Blocksworld hard). In contrast, GPT-4o alone achieves very low performance (less than 0-4%), highlighting its susceptibility to compounding errors in long-horizon compositional reasoning [54, 15]. Our structured visual reasoning approach, incorporating backtracking and verification steps, effectively mitigates these compounding errors, enabling more robust performance even in deep and complex planning problems.
The primary limitation of our method observed in these experiments is the number of incomplete searches (e.g., 40% incomplete in Floor Tiles, 26.67% in Barman) which arise due to the computational budget constraint set. Additionally, we note the method exhibits the highest incorrect rate (60%) on Tetris domain, primarily due to its inherently high branching factor (up to 24 possible actions per state) and complex action parameterization—each action can require up to 7 parameters detailing positions of all sub-tiles, compared to simpler domains like Blocksworld, where actions typically require only 1-2 parameters. Conversely, our largest performance gain over o1-preview occurs in the Barman domain (33.3% vs. 13.3%), likely due to our diagrammatic representation efficiently capturing complex object relationships and high number of interactions, which are challenging to represent for purely textual reasoning models.
@@ 4.4 Ablation Studies
We first assessed the effectiveness of state diagrams by removing them entirely from the inference pipeline (”No Diagram”). This resulted in a substantial performance drop from 90.2% to 58%, highlighting the critical role diagrams play in succinctly representing spatial and relational information. Additionally, the average solution depth of correctly solved instances decreased from 10.07 to 8.28, indicating that without diagrams, the model struggled to solve more complex instances requiring deeper reasoning.
In another experiment, we removed the diagram schema (”No Diagram Schema”) from the child state generation pipeline, directly inferring diagram code from the textual description of the state. This caused performance to drop from 90.2% to 72%. Although less severe than removing diagrams entirely, this drop demonstrates that diagram schemas effectively highlight spatial information from the textual description, aiding accurate diagram generation.
Finally, we tested removing the rendered diagrams, providing only the executable Matplotlib code (”No Code Execution”) when generating the next action. This resulted in the most significant performance drop (90.2% to 24%), even below the GPT-4 baseline (36%), clearly demonstrating that using the diagram code directly distracts the model and impairs reasoning. This also aligns with prior findings that extraneous details negatively impact model performance [47, 34]. These results reinforce the importance of diagrams as compact, intuitive representations that facilitate rapid perceptual inference and clear spatial reasoning [30, 50, 22]. Even though diagrams encode the same information as textual descriptions or code, it is the representation of information in a two-dimensional plane which significantly aids spatial understanding and inference.
@@
We next investigated the effect of branching factor of the inference graph (the number of candidate child states generated per state) on performance. Reducing the branching factor from 4 (our default) to 2 (”2-Branching Factor”) decreased performance from 90.2% to 70%, primarily due to a significant increase in incomplete searches (24% vs. 3.4%). This suggests that the first two candidate states, that in our approach are generated at lower sampling temperatures, include less correct next states, and exploring the third or fourth candidates, generated at higher temperatures, is frequently necessary to find the correct solution path. Thus, reducing branching factor limits randomness and diversity in candidate generation, leading to more incomplete searches. Additionally, the average depth of correctly solved instances decreased from 10.07 to 8.29, indicating difficulty solving problems with deeper solution depths with fewer candidate states. Further reducing the branching factor to 1 (”1-Branching Factor”), effectively converting our graph-based inference into a chain-of-thought approach (with diagrams), caused performance to drop even further to 52%. Despite this significant drop, performance remained substantially higher than the GPT-4 baseline (36%), underscoring the value of diagrams in improving reasoning performance of the LMM even without extensive search.
@@
Finally, we evaluated the importance of our inference optimizations (backtracking and beam search) on top of the graph-of-thought (with diagrams) approach. Removing backtracking (”No Backtracking”) significantly reduced performance from 90.2% to 62%, primarily due to a large increase in incomplete searches (34% vs. 3.4%). This increase arises because the LMM verification steps occasionally produce false negatives, incorrectly invalidating correct states. Without backtracking, the model cannot recover from these errors, leading to incomplete searches as no valid nodes remain at the frontier of the inference graph to expand from. Similarly, removing beam search (”No Beam Search”) reduced performance from 90.2% to 72%, again primarily due to increased incompletes (24% vs. 3.4%). However, unlike backtracking, incompletes here result from exponential growth in the search space, causing the model to exhaust its computational budget (maximum allowed number of states to generate) before reaching the goal state. Indeed, the average number of states generated increased significantly (58.31 vs. 38.89), highlighting beam search’s critical role in efficiently managing combinatorial explosion. Both optimizations proved essential for solving deeper combinatorial problems, as evidenced by the reduced average depth of solved instances (6.61 without beam search, 6.06 without backtracking, vs. 10.07 of our full approach). These results demonstrate that backtracking and beam search are complementary and crucial optimizations for effectively solving complex combinatorial planning problems using graph-based inference.

### v2

@@ 4 Evaluation
To evaluate the proposed approach, we conducted experiments on seven different problem domains, including the popular Blocksworld (simple) Gupta and Nau (1991) and Blocksworld (hard) Valmeekam et al. (2024), as well as 5 new domains contributed in this work, prepared using a popular repository of PDDL problem instance generators ipc . Plan correctness was determined with VAL Howey et al. (2004). Experiments were run on a machine with dual-socket Intel Xeon Gold 5220R processors at 2.2 GHz, 35.75 MB L3, 48 cores per node, 8 nodes total).
Baselines.
We evaluate our approach against both single-inference and search-based inference methods. For single-inference baselines, we evaluate GPT-4o, o1-preview OpenAI (2024c), and o1-mini OpenAI (2024b), each prompted directly with PDDL instances using templates adapted from Valmeekam et al. (2023a). For search-based baselines, we evaluate Graph-of-Thought (GoT), which performs text-only breadth-first search over LLM-generated states, and Optimized GoT, which integrates beam search ($k=4$) into GoT to explore deeper solutions without exhausting the compute budget. We also compare against RAP Hao et al. (2023), which uses Monte Carlo Tree Search (MCTS) with the LLM generating both the world model and the plan. RAP’s compute budget is specified by iteration count; we adopted RAP(20) which was the highest compute budget authors used in their experiments. Although RAP was originally evaluated with LLaMA models, we extended it to use GPT-4o for direct comparison. However, RAP’s implementation requires hard-coded prompts and domain-specific parsers that are only available for Blocksworld, limiting its applicability to the other domains we evaluate. Together, these baselines probe the limits of purely text-based search and single-shot inference and provide strong points of comparison for our diagram-based framework.
Analysis Corrects (Accuracy) Incorrects Incompletes Avg Depth Max Depth Min Depth Avg Num States
@@
Floor Tiles 36% 24% 40% 15.88 25 10 244
Parking 52% 40% 8% 9.6 29 2 149.1
Tetris 38% 60% 2% 6.45 11 4 46.24
Elevator 48% 42% 10% 20.89 26 16 160.30
Barman 30% 40% 30% 24.0 27 22 210.54
@@
Table 2: Analysis of domains. ‘Corrects’: % instances with correct plans; ‘Incorrects’: % with incorrect plans; ‘Incompletes’: % terminated due to state budget; ‘Avg/Max/Min Depth’: avg/max/min number of actions in correct solutions; ‘Avg Num States’: average number of states generated across all instances of the domain.
Evaluated Domains.
We evaluate our method on combinatorial planning problems Ghallab et al. (2004) from the International Planning Competition (IPC) ipc (1998–Present), expressed in PDDL format McDermott et al. (1998). These problems are mostly NP-hard Gupta and Nau (1991). We evaluated our method on seven domains, namely Blocksworld (simple) and Blocksworld (hard) Gupta and Nau (1991); Valmeekam et al. (2024), and 5 other IPC domains: Floor Tiles Helmert (2000), Parking, Tetris Vallati (2015), Elevator, and Barman. For each domain, we generated instances using publicly available instance generators ipc . The Blocksworld domain involves stacking and unstacking blocks. We used the 500 simple instances in PlanBench Valmeekam et al. (2023a) with 3-5 blocks and 50 additional hard instances with 10-20 blocks (same instance complexity as in Valmeekam et al. (2024)). The Floor Tiles domain features robots painting tiles on a grid. We generated 50 PDDL instances for this domain with varying grid sizes (2-3 rows, 3-5 columns) and 1-2 robots. The Parking domain involves rearranging cars in curbs, with 50 instances using 4-5 curbs and 4-6 cars. The Tetris domain requires rearranging Tetris tiles on a grid, with 50 instances using $(4\times 4)$ or $(6\times 6)$ grids. Lastly, the Elevator domain simulates passenger transport in buildings, with 50 instances using 4-5 floors and 10-12 passengers. Refer to Fig. 4 for a sequence of intermediate state diagrams in the correct chain for an example instance of a subset of evaluated domains.
Translating PDDL to Natural Language and Back.
We first translate PDDL domains and instances into natural language. For domain translation, we use a manually engineered five-shot prompt containing translations of five different PDDL domains into natural language. For instance translation, we use a one-shot prompt containing a manually translated instance from the same domain. The pipeline operates on this text. After solving the problem using our pipeline, we translate the resulting natural language action sequence back into PDDL using a one-shot prompt of a random (incorrect) plan with accurate PDDL syntax.
4.1 Results and Analysis
Our main results are summarized in Table 1 which compares the performance of our method against leading models at the time of writing, including the specialized reasoning models o1-preview and o1-mini. Our approach significantly improves the performance of the base GPT-4o model and consistently outperforms strong search-based reasoning baselines across all evaluated domains. On the simpler Blocksworld instances, our method achieves 90.2% accuracy (+54.7 percentage points) and substantially outperforms all baselines, including text-only search-based methods like GoT (50%), Optimized GoT (58%), and RAP (58%), except for a slight disadvantage compared to the leading reasoning model, o1-preview (97.8%). We conjecture two reasons for o1-preview’s advantage: (i) the smaller problem size and shallow solution depth of simpler Blocksworld instances, and (ii) this planning domain is widely used, with many solved instances publicly available online, likely providing extensive exposure to o1 during its training.
On the much harder and deeper Blocksworld (hard) instances and the five new PDDL domains derived from the IPC competition ipc (1998–Present), where new instances are generated and solutions are not publicly available, our method demonstrates superior performance and scalability, supporting the conjectures above. For instance, our approach achieves 78% accuracy on Blocksworld (hard), while o1-preview’s performance drops to 23.65% and RAP’s to just 4%. This trend continues across all new domains, where our method consistently outperforms all baselines, including o1-preview. Standard GoT often fails entirely (e.g., 0% in Floor Tiles and Tetris) due to the combinatorial explosion of states, which exhausts the computational budget. While Optimized GoT improves on this with beam search, its performance still lags significantly behind our visual approach, highlighting that diagrams enable the model to more efficiently capture and analyze complex relational information that is difficult to parse from linear text alone.
Table 2 provides further insights into our method’s performance on these challenging domains. Despite significantly deeper solution paths (with average depths up to 24 actions per instance and up to 244 states generated per instance), our method successfully generates correct plans with as many as 36 sequential actions (Blocksworld (hard)). In contrast, GPT-4o and other search-based methods achieve very low performance (0-4%), highlighting its susceptibility to compounding errors in long-horizon compositional reasoning Valmeekam et al. (2024); Dziri et al. (2023). Our structured visual reasoning approach, incorporating backtracking and verification steps, effectively mitigates these compounding errors, enabling more robust performance even in deep and complex planning problems.
The primary limitation of our method observed in these experiments is the number of incomplete searches (e.g., 40% incomplete in Floor Tiles, 30% in Barman) which arise due to the computational budget constraint. Additionally, we note the method exhibits the highest incorrect rate (60%) on Tetris domain, primarily due to its inherently high branching factor (up to 24 possible actions per state) and complex action parameterization—each action can require up to 7 parameters detailing positions of all sub-tiles, compared to simpler domains like Blocksworld, where actions typically require only 1-2 parameters. Conversely, among the new domains evaluated, our largest performance gain over o1-preview occurs in the Barman domain (30% vs. 10%), likely due to our diagrammatic representation efficiently capturing high number of object interactions, which are challenging to represent for purely textual reasoning models.
Model Generalization.
To assess the generalizability of our framework, we evaluated it on other state-of-the-art LMMs using 50 instances from the Blocksworld (simple) domain. With LLaMA 4, our method increased accuracy from a baseline of 10% (single inference) to 74%, an over 7x improvement in accuracy. The improvement was also observed using Claude 3.5 Sonnet, where accuracy increased from 54.8% (using zero shot single inference as reported in Valmeekam et al. (2024)) to 98%, achieving state-of-the-art performance on this benchmark. These substantial gains demonstrate that the benefits of our approach are not tied to a specific model architecture but stem from the fundamental advantage of using model-generated conceptual diagrams as a reasoning medium.
Runtime and Cost Analysis.
We analyzed the runtime and API costs of our method and all other search-based baselines on 20 instances per domain. On the Blocksworld (simple) domain, our method had a median runtime of 381 ($\sim$6 minutes) seconds and a median cost of \$1.04 per instance. For more complex domains, the median runtime was 1038 seconds ($\sim$17 minutes) with median cost of \$2.98 per instance. Our approach is significantly more efficient than other search-based methods. On average, it was 31% faster and 36% cheaper than the GoT baseline, and 46% faster and 52% cheaper than RAP20+GPT-4o across all domains, while achieving substantially higher accuracy. Compared to the text-only Optimized GoT, incorporating diagrams added 213 seconds in latency and \$0.71 on average, measured across all domains, but this overhead yielded a 30 percentage point accuracy gain.
Ablation Corrects (Accuracy) Incorrects Incompletes Avg Depth Max Depth Avg Num States
GPT-4o + Visual Thinking 90.2% 6.4% 3.4% 10.07 28 38.89
No Diagram (Optimized GoT) 58% 36% 6% 8.28 18 25.86
@@ 4.1 Results and Analysis
Table 3: Ablation Study Results. ‘Corrects’ is our main accuracy metric. See Tab. 2 for columns notation.
@@ 4.2 Ablation Studies
We first assessed the effectiveness of state diagrams by removing them entirely from the inference pipeline ("No Diagram"), resulting in a Graph-of-Thought approach that operates in text-only, rather than a multimodal approach. This led to a substantial performance drop from 90.2% to 58%, highlighting the critical role diagrams play in succinctly representing spatial and relational information. Additionally, the average solution depth of correctly solved instances decreased from 10.07 to 8.28, indicating that without diagrams, the model struggled to solve more complex instances requiring deeper reasoning. In another experiment, we removed the diagram schema ("No Diagram Schema") from the child state generation pipeline, directly inferring diagram code from the textual description of the state. This caused performance to drop from 90.2% to 72%. Although less severe than removing diagrams entirely, this drop demonstrates that diagram schemas effectively capture important spatial information from the textual description, aiding more accurate diagram generation.
Finally, we tested removing the rendered diagrams, providing only the Matplotlib code ("No Code Execution") when generating the next action. This resulted in the most significant performance drop (90.2% to 24%), even below the GPT-4o baseline (35.5%), clearly demonstrating that even though the code encodes the same spatial information as the diagram, the way this information and the interdependencies between objects are represented is crucial for model performance. Using the diagram code directly distracts the model and impairs reasoning, aligning with prior findings that extraneous details negatively impact model performance Shi et al. (2023); Liu et al. (2023b). These results reinforce the importance of diagrams as compact, intuitive representations that facilitate rapid perceptual inference and clear spatial reasoning Larkin and Simon (1987); Tversky (2011); Hegarty (2004), and that it is the representation of information in a multi-dimensional (2D, color-encoded) format that significantly aids spatial understanding and inference in models.
@@
We next investigated the effect of branching factor of the inference graph (the number of candidate child states generated per state) on performance. Reducing the branching factor from 4 (our default) to 2 ("2-Branching Factor") decreased performance from 90.2% to 70%, primarily due to a significant increase in incomplete searches (24% vs. 3.4%). This suggests that exploring the third or fourth candidate child states, generated at higher temperatures, is frequently necessary to find the correct solution path. Thus, reducing branching factor limits randomness and diversity in candidate generation, leading to more incomplete searches. Additionally, the average depth of correctly solved instances decreased from 10.07 to 8.29, indicating difficulty solving problems with deeper solution depths with fewer candidate states. Further reducing the branching factor to 1 ("1-Branching Factor"), effectively converting our graph-based inference into a multimodal chain-of-thought approach (with diagrams), caused performance to drop even further to 52%. Despite this significant drop, performance remained substantially higher than the GPT-4 baseline (36%), underscoring the value of diagrams in improving reasoning performance of the LMM even without extensive search.
@@
Finally, we evaluated the importance of our inference optimizations (backtracking and beam search) on top of the Graph-of-Thought (with diagrams) approach. Removing backtracking ("No Backtracking")—yielding a tree-of-thought method with beam search—reduced performance from 90.2% to 62%, primarily due to a rise in incomplete searches (34% vs. 3.4%). This increase arises because the LMM verification steps occasionally produce false negatives, incorrectly invalidating correct states. Without backtracking, the model cannot recover from these errors, leading to incomplete searches as no validated nodes remain at the frontier of the inference graph to expand from. Similarly, removing beam search ("No Beam Search") dropped performance to 72%, with rate of incomplete searches increasing to 24% . However, unlike backtracking, incompletes here result from exponential growth in the search space, causing the model to exhaust its computational budget (maximum allowed number of states to generate) before reaching the goal state. Indeed, the average number of states generated increased significantly (58.31 vs. 38.89), highlighting beam search’s critical role in managing combinatorial explosion. Both optimizations proved essential for solving deeper combinatorial problems, as shown by the reduced average solution depth without them (6.61 without beam search, 6.06 without backtracking, vs. 10.07 with both). These results demonstrate that backtracking and beam search are complementary and crucial optimizations for effective graph-based combinatorial planning.